OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = 4(one half)x + 1 − 3? Y-intercept of (0, −5), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, −5), starts down on the left, gets closer to y = −3 on the right Y-intercept of (0, 5), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, 5), starts down on the left, gets closer to y = −3 on the right @ganeshie8

ganeshie8 (ganeshie8):

is the given function like this ? \(\large f(x) = 4(\frac{1}{2})^{x + 1} − 3\) ?

OpenStudy (anonymous):

yes @ganeshie8 its like that sorry about that

ganeshie8 (ganeshie8):

then it doesnt match wid any of the options cuz, to get y-intercept, put x = 0 in given function :- \(\large f(x) = 4(\frac{1}{2})^{0 + 1} − 3 \) \(\large f(x) = 4(\frac{1}{2})− 3 = -1 \)

ganeshie8 (ganeshie8):

so y-intercept is (0, -1) you dont have any option wid that :s

ganeshie8 (ganeshie8):

check the function again

OpenStudy (anonymous):

everything is writee and they dont have the option , 0,-1 @ganeshie8

OpenStudy (anonymous):

right*

ganeshie8 (ganeshie8):

for the given function, \((0, -1)\) is the y-intercept. it cannot have any other point as y-intercept lol

ganeshie8 (ganeshie8):

check wid ur teacher if there is a typo in question/options

OpenStudy (anonymous):

lol trust me it isnt its online school a lot of kids takes this

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

that doesnt change the y-intercept :|

OpenStudy (anonymous):

@beccaboo333 do you think the same thing?

OpenStudy (beccaboo333):

give me a moment to loook at this all..

OpenStudy (beccaboo333):

good god, i have no idea. im sorry :(

OpenStudy (beccaboo333):

follow whatever @ganeshie8 says.. he's amazing with math.. literally.. my math savior..

OpenStudy (anonymous):

i know thanks

OpenStudy (dan815):

ganeishi is a good guy

OpenStudy (anonymous):

thanks for telling me dan lol

OpenStudy (dan815):

take a screen shot of your qusetion and send

OpenStudy (anonymous):

i turned in the test and thats the only one i got wrong and im telling you honestly what i pasted here is what it was on the test..

OpenStudy (dan815):